OpenStudy (anonymous):

Solve sin(x)(sinx + 1)=0

OpenStudy (ipwnbunnies):

Break it up into two problems b/c if either one equals 0, the entire thing equals 0. First, sinx = 0. Which angle(s) give us 0? Second, sinx + 1 = 0 ---> sinx = -1. Which angle(s) give us a sin of -1?

OpenStudy (anonymous):

\[a. x=\pm \pi n,x=\pi/2\pm2\pi n\]

OpenStudy (anonymous):

\[b. x=\pi/2\pm2\pi n,x=3\pi/2\pm \pi n\]

OpenStudy (anonymous):

\[c. x=\pm \pi n\]

OpenStudy (anonymous):

\[d. x=\pm \pi n,x=3\pi/2\pm2\pi n\]

OpenStudy (anonymous):

those are the answer options

OpenStudy (ipwnbunnies):

Go through each answer choice and use what I said. You only have to use the first angle given to get your answer.

OpenStudy (anonymous):

so the answer is b

OpenStudy (ipwnbunnies):

No.

OpenStudy (ipwnbunnies):

If it was b, sin(pi/2) = 1, which does not equal zero or -1.